# Provable Sample-Efficient Transfer Learning Conditional Diffusion Models via Representation Learning

### v1

@@ A.2 Proofs of Transfer Learning
$\sqrt{\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}(\widetilde{\ell}(z_{i}^{k},{\bm{f}}_{1},h_{1})-\widetilde{\ell}(z_{i}^{k},{\bm{f}}_{2},h_{2}))^{2}}$ (A.51)
$\qquad\leq 8M^{\frac{1}{2}}\left(\max_{k}\|f_{1}^{k}-f_{2}^{k}\|_{L^{\infty}(\Omega_{R_{1}})}+\gamma_{f}\|h_{1}-h_{2}\|_{L^{\infty}([0,1]^{D_{y}})}\right)+16d_{x}^{\frac{1}{2}}M\exp(-C_{0}^{\prime}R^{2}/2).$
For any $\varepsilon\geq 32d_{x}^{\frac{1}{2}}M\exp(-C_{0}^{\prime}R^{2}/2)$, according to Lemma A.3,
$\log\mathcal{N}(\Phi_{r},\|\cdot\|_{L^{2}(\widehat{\mathbb{P}}_{n}^{(K)})},\varepsilon)$ (A.52)
$\qquad\leq K\log\mathcal{N}(\mathcal{F},\|\cdot\|_{L^{\infty}(\Omega_{R_{1}})},\varepsilon/(16M^{\frac{1}{2}}))+\log\mathcal{N}(\mathcal{H},\|\cdot\|_{L^{\infty}([0,1]^{D_{y}})},\varepsilon/(16\gamma_{f}M^{\frac{1}{2}}))$
@@
Plug in (A.25) and let $\theta=32d_{x}^{\frac{1}{2}}M\exp(-C_{0}^{\prime}R^{2}/2)$,
$\mathcal{R}_{K,n}(\Phi_{r})$ $\leq C_{0}\left(\theta+\int_{\theta}^{2\sqrt{r}}\sqrt{\frac{C_{4}KS_{f}L_{f}\log\left(\frac{L_{f}W_{f}(B_{f}\vee 1)(R\vee T)M}{\varepsilon}\right)+C_{4}S_{h}L_{h}\log\left(\frac{L_{h}W_{h}(B_{h}\vee 1)M\gamma_{f}}{\varepsilon}\right)}{nK}}\mathrm{d}\varepsilon\right)$ (A.53)
@@
$\qquad+C_{0}32d_{x}^{\frac{1}{2}}M\exp(-C_{0}^{\prime}R^{2}/2)$
=:~$=:\widetilde{\mathcal{R}}_{K,n}(r).$
Combine the arguments above, by Lemma A.12 with $B_{0}=8M^{2}\exp(-C_{1}^{\prime}R^{2}),B=4M,b=M$, it holds that with probability no less than $2\displaystyle 1-2nK\exp(-C_{1}^{\prime}R^{2})-\delta/$ 2, for any ${\bm{f}}\in\mathcal{F}^{\otimes K},h\in\mathcal{H}$,
$\mathbb{E}_{z\sim\widehat{\mathbb{P}}^{(K)}}[\widetilde{\ell}(z,{\bm{f}},h)]$ $\leq\frac{2}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\widetilde{\ell}(z_{i}^{k},{\bm{f}},h)+C_{5}M\left(r_{K,n}^{*}+\frac{\log(\log(nK)/\delta)}{nK}\right)$ (A.54)
@@
$\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\widetilde{\ell}(z_{i}^{k},{\bm{f}},h)$ $\leq 2\mathbb{E}_{z\sim\widehat{\mathbb{P}}^{(K)}}[\widetilde{\ell}(z,{\bm{f}},h)]+C_{5}M\left(r_{K,n}^{*}+\frac{\log(\log(nK)/\delta)}{nK}\right)$ (A.55)
@@
where $r_{K,n}^{*}$ is the largest fixed point of ~$\widetilde{\mathcal{R}}_{K,n}$, and it can be bounded by
$r_{K,n}^{*}\leq C_{6}\left(d_{x}^{\frac{1}{2}}M_{f}\exp(-C_{0}^{\prime}R^{2}/2)+\frac{KS_{f}L_{f}\log\left(nL_{f}W_{f}(B_{f}\vee 1)(R\vee T)M\right)+S_{h}L_{h}\log\left(nKL_{h}W_{h}(B_{h}\vee 1)M\gamma_{f}\right)}{nK}\right),$ (A.56)
@@
$\left|\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\ell(x,y,s_{f^{k},h})-\ell(x,y,s_{*}^{k})]-\mathbb{E}_{z\sim\widehat{\mathbb{P}}^{(K)}}[\widetilde{\ell}(z,{\bm{f}},h)]\right|\leq 2M\exp(-C_{1}^{\prime}R^{2}).$ (A.57)
Combine this with (A.54),(A.55),
$\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\ell(x,y,s_{f^{k},h})-\ell(x,y,s_{*}^{\mathbb{P}})]$ $\leq\frac{2}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}[\ell(x_{i}^{k},y_{i}^{k},s_{f^{k},h})-\ell(x_{i}^{k},y_{i}^{k},s_{*}^{k})]$ (A.58)
@@
$\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}[\ell(x_{i}^{k},y_{i}^{k},s_{f^{k},h})-\ell(x_{i}^{k},y_{i}^{k},s_{*}^{k})]$ $\leq\frac{2}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\ell(x,y,s_{f^{k},h})-\ell(x,y,s_{*}^{\mathbb{P}})]$ (A.59)
@@
Plug in the definition of $M=C\left(C_{X}^{{}^{\prime\prime}}R^{6}+M_{f}^{2}+d_{x}\Big{(}\frac{\log(1/T_{0})}{T}+1\Big{)}\right)$ and define $R=C^{\prime}\log^{\frac{1}{2}}(nKd_{x}M_{f}/\delta)$ for some large constant $C^{\prime}$. Hence (A.58) and (A.59) reduce to
$\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\ell(x,y,s_{f^{k},h})-\ell(x,y,s_{*}^{\mathbb{P}})]$ $\leq\frac{2}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}[\ell(x_{i}^{k},y_{i}^{k},s_{f^{k},h})-\ell(x_{i}^{k},y_{i}^{k},s_{*}^{k})]$ (A.60)
@@
$\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}[\ell(x_{i}^{k},y_{i}^{k},s_{f^{k},h})-\ell(x_{i}^{k},y_{i}^{k},s_{*}^{k})]$ $\leq\frac{2}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\ell(x,y,s_{f^{k},h})-\ell(x,y,s_{*}^{\mathbb{P}})]$ (A.61)
@@
$\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\ell^{\mathbb{P}^{k}}(x,y,s_{\widehat{f}^{k},\widehat{h}})]$ (A.62)
@@
$\widehat{{\bm{f}}},\widehat{h}=\operatorname*{arg\,min}_{{\bm{f}}\in\mathcal{F}^{\otimes K},h\in\mathcal{H}}\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\ell(x_{i}^{k},y_{i}^{k},s_{f^{k},h}),$ (A.63)
$\widehat{f}^{\mathbb{P}^{0}}:=\operatorname*{arg\,min}_{f\in\mathcal{F}}\frac{1}{m}\sum_{i=1}^{m}\ell(x_{i}^{0},y_{i}^{0},s_{f,\widehat{h}}).$ (A.64)
@@
$\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}}\mathbb{E}_{(x,y)\sim\mathbb{P}^{0}}[\ell^{\mathbb{P}^{0}}(x,y,s_{\widehat{f}^{\mathbb{P}^{0}},\widehat{h}})]$ $\lesssim\frac{1}{\nu}\inf_{h\in\mathcal{H}}\frac{1}{K}\sum_{k=1}^{K}\inf_{f\in\mathcal{F}}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\ell^{\mathbb{P}^{k}}(x,y,s_{f,h})]+\Delta$ (A.65)
@@
$\log\mathcal{N}_{\mathcal{F}}:=M_{f}^{2}S_{f}L_{f}\log\left(mnL_{f}W_{f}(B_{f}\vee 1)M_{f}T\log(1/\delta)\right),$ (A.66)
$\log\mathcal{N}_{\mathcal{H}}:=S_{h}L_{h}\log\left(nKL_{h}W_{h}(B_{h}\vee 1)M_{f}\gamma_{f}\log(1/\delta)\right).$ (A.67)
@@
$\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}}\mathbb{E}_{(x,y)\sim\mathbb{P}^{0}}[\ell^{\mathbb{P}^{0}}(x,y,s_{\widehat{f}^{\mathbb{P}^{0}},\widehat{h}})]$ (A.68)
@@
Here we apply Proposition A.4 in the first inequality, task diversity in the second inequality, and Proposition A.5 in the fourth. Plug in the definition of $r_{z},r_{x}$ and $\log\mathcal{N}_{\mathcal{F}},\log\mathcal{N}_{\mathcal{H}}$ and we complete the proof. ∎
A.3 Proofs of Meta-Learning
Proposition A.7 (Prop. 3.5).
[…]
$\displaystyle\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathcal{L}(\mathbb{P},h)\leq\frac{2}{K}\sum_{k=1}^{K}\mathcal{L}(\mathbb{P}^{k},h)+C_{P}\left(r_{P}+\frac{\log(1/\delta)}{K}\right),$ (A.69)
$\displaystyle\frac{1}{K}\sum_{k=1}^{K}\mathcal{L}(\mathbb{P}^{k},h)\leq 2\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathcal{L}(\mathbb{P},h)+C_{P}\left(r_{P}+\frac{\log(1/\delta)}{K}\right),$ (A.70)
@@ A.3 Proofs of Meta-Learning
[…]
$\mathcal{R}_{K}(\Phi):=\mathbb{E}_{\bm{\sigma}}\sup_{\varphi\in\Phi}\Big{|}\frac{1}{K}\sum_{k=1}^{K}\sigma_{k}\varphi(\mathbb{P}^{k})\Big{|},\ \bm{\sigma}\sim\text{Unif}(\{-1,1\}^{K}).$ (A.71)
@@
$\Big{\|}\frac{1}{\sqrt{K}}\sum_{k=1}^{K}\sigma_{k}\varphi_{1}(\mathbb{P}^{k})-\frac{1}{\sqrt{K}}\sum_{k=1}^{K}\sigma_{k}\varphi_{2}(\mathbb{P}^{k})\Big{\|}_{\psi_{2}}$ $\leq 4\sqrt{\frac{1}{K}\sum_{k=1}^{K}\|\varphi_{1}(\mathbb{P}^{k})-\varphi_{2}(\mathbb{P}^{k})\|^{2}}$ (A.72)
@@
where $\mathbb{P}_{\text{meta}}^{(K)}:=\frac{1}{K}\sum_{k=1}^{K}\delta_{\mathbb{P}^{k}}$ and $\textbf{diam}\big{(}\Phi_{r},\|\cdot\|_{L^{2}(\mathbb{P}_{\text{meta}}^{(K))}}\big{)}\leq 2\sqrt{r}$. Then by Dudley’s bound [Van Handel, 2014; Wainwright, 2019], there exists an absolute constant $C_{0}$ such that for any $\theta\geq 0$,
$\mathcal{R}_{K}(\Phi_{r})\leq C_{0}\left(\theta+\int_{\theta}^{2\sqrt{r}}\sqrt{\frac{\log\mathcal{N}(\Phi_{r},\|\cdot\|_{L^{2}(\mathbb{P}_{\text{meta}}^{(K)})},\varepsilon)}{K}}\ \mathrm{d}\varepsilon\right).$ (A.73)
@@
$\mathcal{L}(\mathbb{P},h_{1})-\mathcal{L}(P,h_{2})$ $\leq\mathbb{E}_{t,x_{t},y}\left[\Big{|}\|f_{2}(x_{t},h_{1}(y),t)-\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}-\|f_{2}(x_{t},h_{2}(y),t)-\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}\Big{|}\right]$ (A.74)
@@
×$\qquad\qquad\qquad\left.\times\|f_{2}(x_{t},h_{1}(y),t)+f_{2}(x_{t},h_{2}(y),t)-2\nabla_{x}\log p_{t}(x_{t}|y)\|\right]$
@@
In the last inequality we apply $\|f_{i}\|\leq M_{f}$ and $\mathbb{E}_{t,x_{t},y}\|\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}\leq C_{L}$ by Lemma A.9. Moreover,
$\mathbb{E}_{(t,x_{t},y)}\left[\|f_{2}(x_{t},h_{1}(y),t)-f_{2}(x_{t},h_{2}(y),t)\|^{2}\right]$ (A.75)
[…]
@@
$|\mathcal{L}(\mathbb{P},h_{1})-\mathcal{L}(P,h_{2})|\leq C_{3}\left(\gamma_{f}\|h_{1}-h_{2}\|_{L^{\infty}([0,1]^{D_{y}})}+\exp(-C_{1}^{\prime}R_{f}^{2})\right),$ (A.76)
which implies that when $\varepsilon\geq 2C_{3}\exp(-C_{1}^{\prime}R_{f}^{2})$, by Lemma A.3,
$\log\mathcal{N}(\Phi_{r},\|\cdot\|_{L^{2}(\mathbb{P}_{\text{meta}}^{(K)})},\varepsilon)$ $\leq\log\mathcal{N}(\mathcal{H}_{r},\|\cdot\|_{L^{\infty}([0,1]^{D_{y}})},\varepsilon/(2C_{3}\gamma_{f}))$ (A.77)
@@
Plug in (A.73) and let $\theta=2C_{3}\exp(-C_{1}^{\prime}R_{f}^{2})$,
$\mathcal{R}_{K}(\Phi_{r})$ $\leq C_{0}\left(\theta+\int_{\theta}^{2\sqrt{r}}\sqrt{\frac{C_{4}S_{h}L_{h}\log\left(\frac{L_{h}W_{h}(B_{h}\vee 1)C_{3}\gamma_{f}}{\varepsilon}\right)}{K}}\ \mathrm{d}\varepsilon\right)$ (A.78)
@@
=:~$=:\widetilde{\mathcal{R}}_{K}(r).$
According to Lemma A.11 (by setting $B_{0}=0,B=b=C_{L}$), for some absolute constant $C_{5}$, with probability no less than $1-\delta$, we have for any $h\in\mathcal{H}$,
$\displaystyle\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathcal{L}(\mathbb{P},h)\leq\frac{2}{K}\sum_{k=1}^{K}\mathcal{L}(\mathbb{P}^{k},h)+C_{5}C_{L}\left(r_{K}^{*}+\frac{\log(\log(K)/\delta)}{K}\right),$ (A.79)
$\displaystyle\frac{1}{K}\sum_{k=1}^{K}\mathcal{L}(\mathbb{P}^{k},h)\leq 2\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathcal{L}(\mathbb{P},h)+C_{5}C_{L}\left(r_{K}^{*}+\frac{\log(\log(K)/\delta)}{K}\right),$ (A.80)
where $r_{K}^{*}$ is the unique fixed point of ~$\widetilde{\mathcal{R}}_{K}$. And it is easy to show that for some absolute constant $C_{6}$,
$r_{K}^{*}\leq C_{6}\left(C_{3}\exp(-C_{1}^{\prime}R_{f}^{2})+\frac{S_{h}L_{h}\log\left(KL_{h}W_{h}(B_{h}\vee 1)M_{f}\gamma_{f}\right)}{K}\right).$ (A.81)
@@
$\widehat{{\bm{f}}},\widehat{h}=\operatorname*{arg\,min}_{{\bm{f}}\in\mathcal{F}^{\otimes K},h\in\mathcal{H}}\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\ell(x_{i}^{k},y_{i}^{k},s_{f^{k},h}),$ (A.82)
$\widehat{f}^{\mathbb{P}}:=\operatorname*{arg\,min}_{f\in\mathcal{F}}\frac{1}{m}\sum_{i=1}^{m}\ell(x_{i},y_{i},s_{f,\widehat{h}}).$ (A.83)

### v2

@@ B.2 Proofs of Transfer Learning
$\sqrt{\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}(\widetilde{\ell}(z_{i}^{k},{\bm{f}}_{1},h_{1})-\widetilde{\ell}(z_{i}^{k},{\bm{f}}_{2},h_{2}))^{2}}$ (B.51)
[…]
For any $\varepsilon\geq 32d_{x}^{\frac{1}{2}}M\exp(-C_{0}^{\prime}R^{2}/2)$, according to Lemma B.3,
$\log\mathcal{N}(\Phi_{r},\|\cdot\|_{L^{2}(\widehat{\mathbb{P}}_{n}^{(K)})},\varepsilon)$ (B.52)
[…]
@@
Plug in (B.25) and let $\theta=32d_{x}^{\frac{1}{2}}M\exp(-C_{0}^{\prime}R^{2}/2)$,
$\mathcal{R}_{K,n}(\Phi_{r})$ $\leq C_{0}\left(\theta+\int_{\theta}^{2\sqrt{r}}\sqrt{\frac{C_{4}KS_{f}L_{f}\log\left(\frac{L_{f}W_{f}(B_{f}\vee 1)(R\vee T)M}{\varepsilon}\right)+C_{4}S_{h}L_{h}\log\left(\frac{L_{h}W_{h}(B_{h}\vee 1)M\gamma_{f}}{\varepsilon}\right)}{nK}}\mathrm{d}\varepsilon\right)$ (B.53)
@@
[…]
$=:\widetilde{\mathcal{R}}_{K,n}(r).$
Combine the arguments above, by Lemma B.12 with $B_{0}=8M^{2}\exp(-C_{1}^{\prime}R^{2}),B=4M,b=M$, it holds that with probability no less than $1-2nK\exp(-C_{1}^{\prime}R^{2})-\delta/2$, for any ${\bm{f}}\in\mathcal{F}^{\otimes K},h\in\mathcal{H}$,
$\mathbb{E}_{z\sim\widehat{\mathbb{P}}^{(K)}}[\widetilde{\ell}(z,{\bm{f}},h)]$ $\leq\frac{2}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\widetilde{\ell}(z_{i}^{k},{\bm{f}},h)+C_{5}M\left(r_{K,n}^{*}+\frac{\log(\log(nK)/\delta)}{nK}\right)$ (B.54)
@@
$\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\widetilde{\ell}(z_{i}^{k},{\bm{f}},h)$ $\leq 2\mathbb{E}_{z\sim\widehat{\mathbb{P}}^{(K)}}[\widetilde{\ell}(z,{\bm{f}},h)]+C_{5}M\left(r_{K,n}^{*}+\frac{\log(\log(nK)/\delta)}{nK}\right)$ (B.55)
@@
where $r_{K,n}^{*}$ is the largest fixed point of $\widetilde{\mathcal{R}}_{K,n}$, and it can be bounded by
$r_{K,n}^{*}\leq C_{6}\left(d_{x}^{\frac{1}{2}}M_{f}\exp(-C_{0}^{\prime}R^{2}/2)+\frac{KS_{f}L_{f}\log\left(nL_{f}W_{f}(B_{f}\vee 1)(R\vee T)M\right)+S_{h}L_{h}\log\left(nKL_{h}W_{h}(B_{h}\vee 1)M\gamma_{f}\right)}{nK}\right),$ (B.56)
@@
$\left|\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\ell(x,y,s_{f^{k},h})-\ell(x,y,s_{*}^{k})]-\mathbb{E}_{z\sim\widehat{\mathbb{P}}^{(K)}}[\widetilde{\ell}(z,{\bm{f}},h)]\right|\leq 2M\exp(-C_{1}^{\prime}R^{2}).$ (B.57)
Combine this with (B.54),(B.55),
$\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\ell(x,y,s_{f^{k},h})-\ell(x,y,s_{*}^{\mathbb{P}})]$ $\leq\frac{2}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}[\ell(x_{i}^{k},y_{i}^{k},s_{f^{k},h})-\ell(x_{i}^{k},y_{i}^{k},s_{*}^{k})]$ (B.58)
@@
$\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}[\ell(x_{i}^{k},y_{i}^{k},s_{f^{k},h})-\ell(x_{i}^{k},y_{i}^{k},s_{*}^{k})]$ $\leq\frac{2}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\ell(x,y,s_{f^{k},h})-\ell(x,y,s_{*}^{\mathbb{P}})]$ (B.59)
@@
Plug in the definition of $M=C\left(C_{X}^{{}^{\prime\prime}}R^{6}+M_{f}^{2}+d_{x}\Big(\frac{\log(1/T_{0})}{T}+1\Big)\right)$ and define $R=C^{\prime}\log^{\frac{1}{2}}(nKd_{x}M_{f}/\delta)$ for some large constant $C^{\prime}$. Hence (B.58) and (B.59) reduce to
$\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\ell(x,y,s_{f^{k},h})-\ell(x,y,s_{*}^{\mathbb{P}})]$ $\leq\frac{2}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}[\ell(x_{i}^{k},y_{i}^{k},s_{f^{k},h})-\ell(x_{i}^{k},y_{i}^{k},s_{*}^{k})]$ (B.60)
@@
$\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}[\ell(x_{i}^{k},y_{i}^{k},s_{f^{k},h})-\ell(x_{i}^{k},y_{i}^{k},s_{*}^{k})]$ $\leq\frac{2}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\ell(x,y,s_{f^{k},h})-\ell(x,y,s_{*}^{\mathbb{P}})]$ (B.61)
@@
[…]
@@
$\widehat{{\bm{f}}},\widehat{h}=\operatorname*{arg\,min}_{{\bm{f}}\in\mathcal{F}^{\otimes K},h\in\mathcal{H}}\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\ell(x_{i}^{k},y_{i}^{k},s_{f^{k},h}),$ (B.63)
[…]
@@
[…]
@@
$\log\mathcal{N}_{\mathcal{F}}:=M_{f}^{2}S_{f}L_{f}\log\left(mnL_{f}W_{f}(B_{f}\vee 1)M_{f}T\log(1/\delta)\right),$ (B.66)
$\log\mathcal{N}_{\mathcal{H}}:=S_{h}L_{h}\log\left(nKL_{h}W_{h}(B_{h}\vee 1)M_{f}\gamma_{f}\log(1/\delta)\right).$ (B.67)
@@
[…]
@@
Here we apply Proposition B.4 in the first inequality, task diversity in the second inequality, and Proposition B.5 in the fourth. Plug in the definition of $r_{z},r_{x}$ and $\log\mathcal{N}_{\mathcal{F}},\log\mathcal{N}_{\mathcal{H}}$ and we complete the proof. ∎
B.3 Proofs of Meta-Learning
Proposition B.7 (Prop. 3.5).
There exists some constants $C_{1}^{\prime},C_{P}$, such that for $\mathbb{P}^{1},\cdots,\mathbb{P}^{K}\overset{\textit{i.i.d.}}{\sim}\mathbb{P}_{\text{meta}}$, with probability no less than $1-\delta$, we have for any $h\in\mathcal{H}$,
$\displaystyle\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathcal{L}(\mathbb{P},h)\leq\frac{2}{K}\sum_{k=1}^{K}\mathcal{L}(\mathbb{P}^{k},h)+C_{P}\left(r_{P}+\frac{\log(1/\delta)}{K}\right),$ (B.69)
$\displaystyle\frac{1}{K}\sum_{k=1}^{K}\mathcal{L}(\mathbb{P}^{k},h)\leq 2\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathcal{L}(\mathbb{P},h)+C_{P}\left(r_{P}+\frac{\log(1/\delta)}{K}\right),$ (B.70)
@@ B.3 Proofs of Meta-Learning
Given $\mathbb{P}^{1},\cdots,\mathbb{P}^{K}\overset{\textit{i.i.d.}}{\sim}\mathbb{P}_{\text{meta}}$, we define the empirical Rademacher complexity of a function class $\Phi$ defined on the set of distribution $\mathcal{P}(\mathbb{R}^{d_{x}}\times[0,1]^{D_{y}})$ as
$\mathcal{R}_{K}(\Phi):=\mathbb{E}_{\bm{\sigma}}\sup_{\varphi\in\Phi}\Big|\frac{1}{K}\sum_{k=1}^{K}\sigma_{k}\varphi(\mathbb{P}^{k})\Big|,\ \bm{\sigma}\sim\text{Unif}(\{-1,1\}^{K}).$ (B.71)
@@
$\Big\|\frac{1}{\sqrt{K}}\sum_{k=1}^{K}\sigma_{k}\varphi_{1}(\mathbb{P}^{k})-\frac{1}{\sqrt{K}}\sum_{k=1}^{K}\sigma_{k}\varphi_{2}(\mathbb{P}^{k})\Big\|_{\psi_{2}}$ $\leq 4\sqrt{\frac{1}{K}\sum_{k=1}^{K}\|\varphi_{1}(\mathbb{P}^{k})-\varphi_{2}(\mathbb{P}^{k})\|^{2}}$ (B.72)
@@
where $\mathbb{P}_{\text{meta}}^{(K)}:=\frac{1}{K}\sum_{k=1}^{K}\delta_{\mathbb{P}^{k}}$ and $\textbf{diam}\big(\Phi_{r},\|\cdot\|_{L^{2}(\mathbb{P}_{\text{meta}}^{(K))}}\big)\leq 2\sqrt{r}$. Then by Dudley’s bound [Van Handel, 2014, Wainwright, 2019], there exists an absolute constant $C_{0}$ such that for any $\theta\geq 0$,
$\mathcal{R}_{K}(\Phi_{r})\leq C_{0}\left(\theta+\int_{\theta}^{2\sqrt{r}}\sqrt{\frac{\log\mathcal{N}(\Phi_{r},\|\cdot\|_{L^{2}(\mathbb{P}_{\text{meta}}^{(K)})},\varepsilon)}{K}}\ \mathrm{d}\varepsilon\right).$ (B.73)
@@
$\mathcal{L}(\mathbb{P},h_{1})-\mathcal{L}(P,h_{2})$ $\leq\mathbb{E}_{t,x_{t},y}\left[\Big|\|f_{2}(x_{t},h_{1}(y),t)-\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}-\|f_{2}(x_{t},h_{2}(y),t)-\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}\Big|\right]$ (B.74)
@@
$\qquad\qquad\qquad\left.\times\|f_{2}(x_{t},h_{1}(y),t)+f_{2}(x_{t},h_{2}(y),t)-2\nabla_{x}\log p_{t}(x_{t}|y)\|\right]$
@@
In the last inequality we apply $\|f_{i}\|\leq M_{f}$ and $\mathbb{E}_{t,x_{t},y}\|\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}\leq C_{L}$ by Lemma B.9. Moreover,
$\mathbb{E}_{(t,x_{t},y)}\left[\|f_{2}(x_{t},h_{1}(y),t)-f_{2}(x_{t},h_{2}(y),t)\|^{2}\right]$ (B.75)
$\leq\mathbb{E}_{t,y}\left[\int\|f_{2}(x_{t},h_{1}(y),t)-f_{2}(x_{t},h_{2}(y),t)\|^{2}p_{t}(x_{t}|y)dx_{t}\right]$
$\leq\mathbb{E}_{t,y}\left[\int_{\|x_{t}\|_{\infty}\leq R_{f}}\|f_{2}(x_{t},h_{1}(y),t)-f_{2}(x_{t},h_{2}(y),t)\|^{2}p_{t}(x_{t}|y)dx_{t}+4M_{f}^{2}\mathbb{P}(\|x_{t}\|_{\infty}>R_{f}|y)\right]$
@@
$|\mathcal{L}(\mathbb{P},h_{1})-\mathcal{L}(P,h_{2})|\leq C_{3}\left(\gamma_{f}\|h_{1}-h_{2}\|_{L^{\infty}([0,1]^{D_{y}})}+\exp(-C_{1}^{\prime}R_{f}^{2})\right),$ (B.76)
which implies that when $\varepsilon\geq 2C_{3}\exp(-C_{1}^{\prime}R_{f}^{2})$, by Lemma B.3,
$\log\mathcal{N}(\Phi_{r},\|\cdot\|_{L^{2}(\mathbb{P}_{\text{meta}}^{(K)})},\varepsilon)$ $\leq\log\mathcal{N}(\mathcal{H}_{r},\|\cdot\|_{L^{\infty}([0,1]^{D_{y}})},\varepsilon/(2C_{3}\gamma_{f}))$ (B.77)
@@
Plug in (B.73) and let $\theta=2C_{3}\exp(-C_{1}^{\prime}R_{f}^{2})$,
$\mathcal{R}_{K}(\Phi_{r})$ $\leq C_{0}\left(\theta+\int_{\theta}^{2\sqrt{r}}\sqrt{\frac{C_{4}S_{h}L_{h}\log\left(\frac{L_{h}W_{h}(B_{h}\vee 1)C_{3}\gamma_{f}}{\varepsilon}\right)}{K}}\ \mathrm{d}\varepsilon\right)$ (B.78)
@@
$=:\widetilde{\mathcal{R}}_{K}(r).$
According to Lemma B.11 (by setting $B_{0}=0,B=b=C_{L}$), for some absolute constant $C_{5}$, with probability no less than $1-\delta$, we have for any $h\in\mathcal{H}$,
$\displaystyle\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathcal{L}(\mathbb{P},h)\leq\frac{2}{K}\sum_{k=1}^{K}\mathcal{L}(\mathbb{P}^{k},h)+C_{5}C_{L}\left(r_{K}^{*}+\frac{\log(\log(K)/\delta)}{K}\right),$ (B.79)
$\displaystyle\frac{1}{K}\sum_{k=1}^{K}\mathcal{L}(\mathbb{P}^{k},h)\leq 2\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathcal{L}(\mathbb{P},h)+C_{5}C_{L}\left(r_{K}^{*}+\frac{\log(\log(K)/\delta)}{K}\right),$ (B.80)
where $r_{K}^{*}$ is the unique fixed point of $\widetilde{\mathcal{R}}_{K}$. And it is easy to show that for some absolute constant $C_{6}$,
$r_{K}^{*}\leq C_{6}\left(C_{3}\exp(-C_{1}^{\prime}R_{f}^{2})+\frac{S_{h}L_{h}\log\left(KL_{h}W_{h}(B_{h}\vee 1)M_{f}\gamma_{f}\right)}{K}\right).$ (B.81)
@@
$\widehat{{\bm{f}}},\widehat{h}=\operatorname*{arg\,min}_{{\bm{f}}\in\mathcal{F}^{\otimes K},h\in\mathcal{H}}\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\ell(x_{i}^{k},y_{i}^{k},s_{f^{k},h}),$ (B.82)
[…]